find two consecutive integers whose product is 42
Answer by richwmiller(17219)

(n+1)*n=42
n^2+n=42
n^2+n-42=0
we want factor of 42 wuth a difference of one
7 and 6 come to mind
1,42
2,21
3,14
(n-6)(n+7)=0
The problem doesn't say positive integers so we have two answers
n=6 and n=-7
6 and 7
and
-7 and -6